OpenStudy (anonymous):

ab + cd = 5 Make d the subject

OpenStudy (anonymous):

ab + cd = 5 cd=5-ab d=(5-ab)/c

OpenStudy (anonymous):

cd = -ab + 5 -ab + 5 d = -------- c

OpenStudy (anonymous):

which one is correct?

OpenStudy (anonymous):

-ab+5 or 5-ab?

OpenStudy (anonymous):

both mean the same thing...... just their places have been interchanged.... normally in algebra we write the constant last so i hv shifted 5 to the end and -ab to the front........

OpenStudy (anonymous):

\[d=(5-ab)/c\]